### v1

# HateDebias: On the Diversity and Variability of Hate Speech Debiasing

Abstract
Hate speech on social media is ubiquitous but urgently controlled. Without detecting and mitigating the biases brought by hate speech, different types of ethical problems. While a number of datasets have been proposed to address the problem of hate speech detection, these datasets seldom consider the diversity and variability of bias, making it far from real-world scenarios. To fill this gap, we propose a benchmark, named HateDebias, to analyze the model ability of hate speech detection under continuous, changing environments. Specifically, to meet the diversity of biases, we collect existing hate speech detection datasets with different types of biases. To further meet the variability (i.e., the changing of bias attributes in datasets), we reorganize datasets to follow the continuous learning setting. We evaluate the detection accuracy of models trained on the datasets with a single type of bias with the performance on the HateDebias, where a significant performance drop is observed. To provide a potential direction for debiasing, we further propose a debiasing framework based on continuous learning and bias information regularization, as well as the memory replay strategies to ensure the debiasing ability of the model. Experiment results on the proposed benchmark show that the aforementioned method can improve several baselines with a distinguished margin, highlighting its effectiveness in real-world applications.
1 Introduction
Hate speech [1, 2] is frequent on social media, which has become a critical social problem and is desired to solve. One mainstream solution is to employ the technique of machine learning to detect and further mitigate hate speech [3, 4, 5], hence a safe and inclusive environment of the internet can be protected by removing these hate speeches. To well devise algorithms for hate speech detection, one of the most important steps is to provide benchmarks that are close to real-world scenarios.
There are different types of benchmarks to evaluate the hate speech detection algorithm. Previously, ETHOS [6] is built from YouTube and Reddit comments validated through a crowdsourcing platform, which is designed to recognize the target of hate speech. The Measuring Hate Speech corpus [7] is devised to identify the target community and classify a sentence is a hate speech or not. In summary, these datasets are devised to recognize the target of hate speech. Recently, researchers have aimed to address the problem of debiasing and propose different datasets. Specifically, RaceBiasHate [8] uses the Twitter data annotated for hate speech and abusive language to address the racial bias. Geo-CulturalBiasHate [9] evaluates cultural bias in hate speech datasets by leveraging two interrelated cultural proxies: language and geography. Furthermore, the MultilingualBiasHate corpus [10] further considers different types of hate speech like age, country, gender, and race. In summary, existing datasets usually focus on the target of hate speech and seldom consider different types of biases.
Figure 1: Comparison of existing datasets and our dataset with real-world scenarios. Figure (a) shows that existing datasets typically contain a single type of bias, which may lead to suboptimal performance of models when dealing with real-world scenarios involving multiple and varying biases. Figure (b) indicates that in real-world scenarios, there are multi-types and continuous-varying biases. Figure (c) demonstrates our constructed HateDebias dataset, which simulates the continuous emergence of various types of hate speech biases, making the dataset distribution more similar to real-world scenarios and thereby improving model performance in practical applications.
Considering the multiple types of biases of hate speech is essential to developing the hate speech debiasing models, as it prevents models from ignoring different biases during training. Although these datasets have boosted the development of hate speech detection algorithms to some extent, several algorithms trained on these datasets can hardly achieve ideal performance in practice since they are still far from real-world scenarios. As shown in Figure 1 (a), since existing datasets contain a single type of bias, the models trained on these datasets may achieve suboptimal performance on real-world scenarios with multiple and time-varying biases in Figure 1 (b). To address this challenge, it is desired to build a benchmark that close to real-world application. As shown in Figure 1 (c), when simulating how the multiple types of hate speech occur continuously, we can build a dataset with a similar distribution as the real-world scenarios, so the models may achieve ideal performance.
Table 1: Comparisons of different datasets. BI: Biased Information; D: Diversity; V: Variability.
Dataset BI D V
ETHOS [6]
Measuring Hate Speech [7]
Geo-CulturalBiasHate [9]
RaceBiasHate [8]
MultilingualBiasHate [10]
[…]
To achieve this, we first formalize the hate speech scenarios with multiple types and dynamically varying biases as continuous debiasing tasks. Our goal is to de-bias each sub-dataset with different types of hate speech biases, and meanwhile, to maintain high performance on downstream tasks. Sequentially, we construct a real-world dataset, HateDebias, to serve as a benchmark for these continuous debiasing tasks. It consists of 23276 hate speech text with 4 types of bias attributes, including age, country, gender, and ethnicity. Specifically, we simulate the emergence of various new biases at different stages based on the [11] and [10], such that each type of bias corresponds to a specific sub-dataset in HateDebias. As shown in Table 1, one can see that our dataset offers greater diversity and variability compared to existing hate speech detection datasets. Moreover, we propose a continual-learning-based framework, leveraging a bias information regularization (BIR) strategy and a memory replay (MR) strategy to better understand and tackle biases in the training process.
Contribution.
(1) We explicitly formalize the continuous debiasing task that tackles the continuous debiasing scenario with multi-types and continuous-varying bias. (2) We construct a real-world dataset, called HateDebias, where any sequence of sub-dataset in HateDebias corresponds to continuous-varying bias. (3) We propose a simple yet effective framework for continuous debiasing tasks, which provides a potential direction to improve continuous debiasing tasks.
Figure 2: Overview of continuous debiasing task. The proposed HateDebias benchmark includes three aspects: task definition, dataset construction and practical framework.
2 Benchmark Continuous Debiasing Tasks
In this section, we aim to benchmark the continuous debiasing task. As shown in Figure 2, the proposed HateDebias benchmark includes three aspects: task definition, dataset construction and practical framework. Specifically, we first explicitly define the continuous debiasing task that aims to continually update the model to adapt to new biases as they emerge (Sec. 2.1). Secondly, we propose a real-world dataset in which various biases are introduced periodically, used as the benchmark for evaluating continuous debiasing tasks (Sec. 2.2). Finally, we design a continual-learning-based framework incorporating techniques like bias regularization and memory replay to effectively debias over time, as a general solution for handling continuous debiasing tasks (Sec. 2.3).
2.1 Continuous Debiasing Task Definition
We focus on attribute-incremental continual learning scenarios. For example, consider a hate speech detection application used by a social media platform. Initially, the model is deployed to detect hate speech and mitigate racial bias in user-generated content. This model, while effective, must operate in a dynamic environment where new types of biases would emerge. Suppose the platform then observes an increasing trend of gender-biased hate speech. When this new bias is detected, the existing model must be adapted to address and mitigate this bias without neglecting its original capability.
Let $D_{i}$ be a dataset and $A_{i}$ be the bias attribute corresponding to $D_{i}$, denoted their relationship as $BiasArr(D_{i})=A_{i}$, the continuous debias task is defined as follows:
Definition 2.1 (Continuous Debiasing Task).
Given a hate speech dataset $\mathcal{D}=\{D_{1},\cdots,D_{n}\}$ with its corresponding set of bias attribute $\mathcal{A}=\{A_{1},\cdots,A_{n}\}$ that satisfies $BiasArr(D_{i})=A_{i}$, the continuous debiasing task, denote as a task sequence $\mathcal{T}_{\textit{de-bias}}=\{\operatorname{T_{1}},\cdots,\operatorname{T_{n}}\}$, aims to de-bias each sub-dataset sequentially, i.e., $\operatorname{MBias}(\operatorname{T_{i}}(D_{i},A_{i}))\neq A_{i}$, $i=1,2,\cdots,n$, where $\operatorname{MBias}(\cdot)$ is a function that reports the type of hate speech bias of a model, $\operatorname{T_{i}}$ is a sub-task that accept dataset $D_{i}$ and its corresponding bias attribute $A_{i}$ as input, and output an adjusted model.
It is worth noting that the order of bias attribute in $\mathcal{A}$ can be arbitrary, implying the diversity and variability of bias attribute for the dataset $\mathcal{D}$. Moreover, the goal of our work is not only to debias the hate speech bias embedded in the model but also to maintain performance in downstream tasks applied to the dataset $\mathcal{D}$, such as classification tasks.
There are two key characteristics of the continuous debiasing task: diversity and variability. We will properly define them to induce the construction of the benchmark dataset. Diversity aims to emphasize the multi-type bias attributes within the dataset $\mathcal{D}$, e.g., $|\mathcal{A}|\geq 3$ where ||$\cdot$|| ⋅ | represents the dimension. On the other hand, variability represents the difference in bias attributes across various stages of the task. For instance, in a continuous debiasing task sequence $\mathcal{T}_{\textit{de-bias}}=\{\operatorname{T}_{1},\cdots,\operatorname{T}_{n}\}$, $\operatorname{T}_{i}$ aims to debias the bias attribute $A_{i}$ while $\operatorname{T}_{i+1}$ addresses the bias attribute $A_{j}$, with ${A_{i}}\neq A_{j}$.
2.2 Continuous Debiasing Dataset
After a well-defined continuous debiasing task, this section will provide the detailed construction of our benchmark dataset, HateDebias, which can be used to evaluate the continuous debiasing task and accordingly encourage the development of related methods later. To better simulate scenarios that are close to the real-world environment, the HateDebias dataset is required to satisfy two conditions in our construction procedure: diversity and variability of biases.
Diversity.
There are some hate speech datasets with various bias attributes. Huang et al. [10] gather and annotate a Twitter dataset with four types of demographic attributes, including age, country, gender, and ethnicity. Through the utilization of the Twitter streaming API [^1] https://developer.twitter.com/, they obtain tweets that were annotated with hate speech labels [12, 13], along with the corresponding user profiles. The age, race, and gender attributes of the users are inferred from analyzing their profile images through Face++ [^2] https://www.faceplusplus.com/, a computer vision API capable of estimating demographic traits. The country attribute of each user is obtained by extracting geographic information and using the Google Maps API [^3] https://maps.googleapis.com to acquire structured location data. All attributes and the hate label are converted into binary format. Thus, we select the English text dataset from the hate speech detection task corpus from the Twitter platform Huang et al. [10]. The diversity of bias attributes is guaranteed since this corpus is annotated with four types of demographic attributes, including age, country, gender, and ethnicity. In this dataset, the validation set consists of 12681 samples, and the test set comprises 12682 samples, following the dataset split by Huang et al. [10].
Table 2: Label and four attributes distribution of the training set. The row represents datasets and the column shows the attributes and labels within these datasets, and the table values detail the distribution of each attribute and label. The term “Dataset” refers to the aggregate of all four subsets.
Attribute Value Dataset Subset
Age Gender Country Ethnicity
Age elder 10561 2640 2639 2643 2639
median 12715 3179 3180 3176 3180
Gender male 8850 2212 2213 2212 2213
female 14426 3607 3606 3607 3606
Country non-US 7418 1855 1854 1854 1855
US 15858 3964 3965 3965 3964
Ethnicity non-white 15480 3871 3871 3869 3869
white 7796 1948 1948 1950 1950
Label non-hate 17026 4255 4258 4256 4257
hate 6250 1564 1561 1563 1562
Variability.
To achieve the variability of bias types in dataset $\mathcal{D}$, our strategy is to generate sub-datasets with various types of bias and rearrange the sequence of these sub-datasets. We first collect a training set of 23,276 samples from the original training set, which consists of 59,179 samples, by removing samples with missing values, specifically those missing demographic attributes. This is because the continuous debiasing task requires the known bias attribute for its corresponding dataset. Based on the collected dataset with four bias attributes, we further partition it into four distinct subsets. To ensure distribution consistent with the collected data, the key issue in this partitioning is maintaining the same proportion of samples in each subset as in the overall dataset [14]. We address this issue using an equal proportion division method. Specifically, we randomly stratify the entire dataset based on a combination of labels (hate speech and non-hate speech) and demographic attributes (age, gender, country, and ethnicity), constraining (i) the proportion of each attribute in various sub-dataset is the same as the completed data, and (ii) there are not overlap sample between different sub-dataset. The results are shown in Table 2, there are four subsets, ‘Age’, ‘Gender’, ‘Country’, and ‘Ethnicity’. One can see that the proportion of each attribute among them is consistent with the completed data. To simulate the scenarios with continuous and changing bias types, we first assume only one explicit bias attribute in the sub-dataset and no other bias attribute can be observed. It means that, for any subset $D_{i}$, only the bias attribute $A_{i}$ is under-investigated. Furthermore, we randomly generate the sequence of these sub-datasets such that for the completed dataset $\mathcal{D}$ with its partition $\{D_{1},D_{2},D_{3},D_{4}\}$, there exists a significant changing of bias attribute.
Example.
We provide sample examples from the HateDebias Dataset in Table  3, with one sample per row. We randomly selected four examples from subsets containing 5819 samples, covering both biased and unbiased cases for each sensitive attribute. Each color sequence represents specific bias variations in the HateDebias dataset, reflecting real-world biases. The dataset includes various sensitive attributes, such as age, country, gender, and ethnicity, converted into binary format. Table  2 details the number of samples and values for each attribute. The text field contains examples from a Twitter dataset for hate speech classification, with sensitive information masked to protect privacy. The symbol ✔ indicates bias for the corresponding attribute, while ✘ indicates no bias.
Table 3: Examples of the HateDebias dataset.
Sensitive
Attribute
Value Text Bias
elder
user user royal calcutta was established in1829 and isnt in top 10 oldest golf clubs. it is the oldest
outside britain tho.
elder and then there was one hashtag i think the old man carl theory is pretty sick all he has left is the magnum…
median user and it reminds me too much of parties i had with friends when i was younger . minus the cam, lol.
Age median "ah feminism is still going strong, " "at least weŕe still two young, hot, blondes." " hashtag"
non-US how one woman used hashtag to reclaim her muslim american identity url via user url
non-US user user the only people that are stupid is you, vanderbilt did not build america, he owned or poorly paid…
US user user that’s only because all of the african american population came out to vote. you really do…
Country US
latinos who think they’re woke by choosing coke over pepsi and not knowing how bad coke has screwed latin
america hashtag
non-white what if my skin is so white it glows, and i’ve got blue hair? hashtag url
non-white
user there is literally nothing more creatively pathetic than taking an existing film and just making a
black or female version.
white
cause it don’t pay to inform your own, it’s more profitable to point the finger at the white man all the time
and not take any ownership.
Ethnicity white
user i haven’t seen anything other than clips and then read a bunch of reviews from black people about
why they hated it
male user user user user from left to right, it’s like the steps of male pattern hair loss.
male
hashtag bunch of 1st world feminists dictating to a third one country in sexist
racism way how a male positive ht is bad for women
female
how many top grossing films with a strong female lead do we have to have before hollywood acknowledges
a winning strategy ?
Gender female rt user : user 2006 self made man: norah vincent chooses female privilege over male privilege url
2.3 Continual-learning-based Framework
Figure 3: Continual-learning-based Framework. Memory replay selects samples from prior tasks to solidify previously acquired knowledge. Bias information regularization maintains shared knowledge across tasks and integrate bias-specific information.
To further understand the continuous debiase task and verify the benchmark dataset, we develop a continual-learning-based framework (CLF) for continuous debiasing, aimed at improving model performance and fairness by learning attribute-irrelevant features to reduce bias. Our method uses a two-stage training process: the initial training stage and the memory replay stage. Initially, the model trains on current task data to acquire new knowledge. In the memory replay stage, representative samples from previous tasks are selected to update the memory for consolidating prior knowledge. To preserve shared knowledge across tasks while integrating bias-specific knowledge, we propose a bias information regularization technique to separate bias information into a generic space and a bias-specific space, facilitating interaction and helping the model adapt to various debiasing tasks. The more detailed implementation of our framework is provided in Appendix A4.
2.3.1 Bias Information Regularization
Regularization technology separates generic and bias-specific information in hidden representations, facilitating sentence structure and context analysis. For a sentence $x$, a pre-trained language model $B(\cdot)$ produces hidden representations $h$, containing generic and bias-specific information. We then construct two networks, $G(\cdot)$ and $S(\cdot)$, to extract generic ($g$) and bias-specific ($s$) representations from hidden representations $h$. For bias-specific representation $s$, we introduce a bias-identifier prediction task that classifies examples based on their bias type using the bias-specific representation $s$. The loss function for the bias-identifier predictor is $\mathcal{L}_{task}$ (see Sec. A4.2 for detailed description).
To address catastrophic forgetting in debiasing pre-trained language models via continual learning, we propose a representation regularization approach for generic $g$ and bias-specific $s$ representations. This approach separately preserves knowledge in $g$ and $s$, promoting positive transfer and reducing forgetting. Before training on task $t$, we compute these representations for all sentences in the training set $D_{t}$ and memory buffer $M_{t}$. Using the encoder $B^{t-1}(\cdot)$ and networks $G^{t-1}(\cdot)$, $S^{t-1}(\cdot)$ from task $t-1$, we calculate $G^{t-1}(B^{t-1}(x))$ and $S^{t-1}(B^{t-1}(x))$. These are combined with task $t$’s representations for L2 norm regularization, constraining changes in both representation spaces. Thus, during the learning in debiasing task $t$, we impose two regularization losses respectively:
$\mathcal{L}_{g}=E_{x\in D_{t}\cup M_{t}}||G^{t-1}(B^{t-1}(x))-G^{t}(B^{t}(x))||_{2},$ (1)
$\mathcal{L}_{s}=E_{x\in D_{t}\cup M_{t}}||S^{t-1}(B^{t-1}(x))-S^{t}(B^{t}(x))||_{2},$ (2)
where $||\cdot||_{2}$ is the L2 norm regularization operation.
2.3.2 Memory Replay
To efficiently use the memory buffer $M$ while maintaining balanced replay, we selectively store a small number of samples from the previous debiasing task’s training set. We explore strategies to recall specific samples to mitigate model bias, which arises from performance variations across groups defined by a sensitive attribute. Replaying low-accuracy samples, particularly those associated with a sensitive attribute, can effectively reduce bias. In the $t$-th debiasing stage, before fine-tuning and debiasing the model, we select the corpus $D_{r}$ for stage $r(r\textless t)$. First, the class predictor $f_{cls}(\cdot)$ is used to predict $D_{r}$, and the probability list $P$ corresponding to the true label of each sample is obtained. We set the number of retained samples as $\gamma|D_{r}|$, and obtain the corresponding probability threshold as $\beta$, then the samples retained in stage $r$ are:
$M_{t}^{r}=\{x\in D_{r}|P(x)\textless\beta\}(M_{t}^{r}\subseteq M_{t}).$ (3)
2.3.3 Overall Objective
When performing the $t$-th debiasing task, we simultaneously leverage the generic representation $g$ and the bias-specific representation $s$ for text classification. To optimize model performance, we minimize the cross-entropy loss $\mathcal{L}_{cls}$ (see Sec. A4.3 for detailed description). In addition, for a given sentence $x$, we adopt the given debiaser $H_{t}(\cdot)$ to mitigate attribute $A_{t}$ in the text. The debiaser $H_{t}(\cdot)$ is designed to be highly flexible and extensible, allowing for modifications as needed. In our experiments, we utilize four debiasing methods as the backbones of the debiaser $H_{t}(\cdot)$, namely FGM [15], PGD [16], ATC [17, 18, 19] and CL [20, 21, 22]. We define the debiaser’s loss as $\mathcal{L}_{d}$ (see Sec. A4.4 for detailed description). The final loss of our continual debiasing model is:
$\mathcal{L}=\mathcal{L}_{cls}+\alpha\cdot\mathcal{L}_{d}+\sigma\cdot(\mathcal{L}_{g}+\mathcal{L}_{s}+\mathcal{L}_{task}).$ (4)
We set the coefficient of the first loss terms to 1 for simplicity and only introduce two loss weights to tune: $\sigma$ and $\alpha$. The loss weight $\alpha$ represents the debiasing degree of the model, while the loss weight $\sigma$ controls the regularization degree of bias information.
3 Experiment
We design multiple sets of experiments to validate the effectiveness of our proposed model on the proposed dataset. First, we evaluate that existing methods can hardly achieve the ideal performance in the multiple biases while the proposed method can solve this problem, which is shown in 3.2 and Sec. 3.3. Sequentially, to further evaluate that our method can achieve stable performance under the continuous scenarios of our datasets, we further consider different numbers of biases in the continuous scenarios, which is shown in Sec. 3.4. Moreover, we evaluate that our method can avoid catastrophic forgetting as shown in Sec. 3.5. Besides, we conduct ablation experiments, parameter search experiments, statistical significance tests, explore the effects of different sequence orders, and different pre-trained language models. Please refer to more experiment results in Appendix A7.
3.1 Setups
Baselines
We consider seven baselines: Fine-tune [23], Experience Replay (ER) [24, 11], IDBR [25], Multi-task Learning (MTL) [26], Two-level Knowledge Distillation (TKD) [27], Representation Projection Invariance (REPINA) [28] and (Instance-wise Relation Distillation (IRD) [29] to combine debiasing methods to verify the effectiveness of various strategies and our proposed method. Given that our experimental setting involves attribute-incremental continual learning, we transform the attribute-incremental setting into a debiasing-task-incremental setting.
Model Details
Table 4: The optimal hyperparameter values of each debiasing method.
Method $\gamma$ $\alpha$ $\sigma$
FGM 0.1 1 0.1
PGD 0.1 1 0.05
ATC 0.05 0.1 0.05
CL 0.1 0.1 0.1
All experiments are carried out based on PyTorch and an A30 with 24 GB of memory. We design our framework based on Transformers. Furthermore, in terms of the pre-trained language model, we choose the bert-base-uncased model [30] as the backbone of our framework. We set the same hyper-parameters with a fixed initialization seed for our model’s training, where the batch size is 32 and the feature dimension is 768. The max length of the input samples is 32. The BERT encoder with a learning rate of 5e-5 is optimized utilizing Adam. For the MTL model, we set the training epoch to 10, while for other continual learning methods, we set the number of epochs to 5 for each stage. We perform a grid search on the hyperparameters in the model. The optimal hyperparameter values of each debiasing method are shown in Table 4. In order to avoid the contingency caused by the order of debiased attributes, we conduct experiments with $A_{4}^{4}=24$ combinations of 4 attributes.
Evaluation Metrics
Formally speaking, in the continuous debiasing task, we require that a certain type of bias cannot be detected in the model after the debiasing function $MBais(\cdot)$. Unlike traditional task-incremental continual learning tasks, we perform attribute-incremental continual learning on the same downstream task. We adopt the accuracy (Acc) and macro-average $F_{1}$ (F1-macro) value to evaluate the performance of models in the downstream task. To evaluate group fairness, we measure the equality differences of false positive rates (FPR) for each bias. In terms of the overall fairness of the model, the average false positive rate of four biases is calculated to measure, which is defined as the averaged attribute bias (AAB). Since there are 24 sequence settings in the attribute-incremental continual learning task, for continual-learning-based methods, we calculate the average of these metrics for all task sequences. For each metric, we also calculate the variance.
Table 5: Main results. The columns and rows are the evaluation metrics and the names of each method respectively. Values in brackets are the variances of the metrics. For the static debiasing method, since there are no multiple sequence experiments, we denote the variance as 0.00. MTL represents a static debiasing approach that addresses multiple sensitive attributes concurrently. Fine-tune constitutes the most straightforward experimental setup for continual debiasing tasks. Furthermore, ER, IDBR, REPINA, IRD and TKD are prevalent methods in continuous learning for continual debiasing tasks.
Method Debiasing Method Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ FPR(Gender) $\downarrow$ FPR(Age) $\downarrow$ FPR(Country) $\downarrow$ FPR(Ethnicity) $\downarrow$ AAB $\downarrow$
MTL [26] No-debias 85.90(0.00) 83.00(0.00) 4.77(0.00) 7.77(0.00) 5.35(0.00) 4.43(0.00) 5.58(0.00)
PGD 86.79(0.00) 83.59(0.00) 3.05(0.00) 7.37(0.00) 3.21(0.00) 3.87(0.00) 4.37(0.00)
FGM 87.19(0.00) 83.99(0.00) 3.92(0.00) 6.05(0.00) 4.97(0.00) 4.18(0.00) 4.78(0.00)
ATC 85.48(0.00) 82.32(0.00) 6.11(0.00) 8.28(0.00) 5.55(0.00) 3.87(0.00) 5.95(0.00)
CL 88.12(0.00) 84.38(0.00) 4.76(0.00) 7.04(0.00) 6.26(0.00) 5.17(0.00) 5.81(0.00)
Fine-tune [23] No-debias 86.73(2.92) 83.51(3.23) 5.55(0.91) 8.67(2.16) 5.97(1.67) 4.51(1.10) 6.17(0.65)
PGD 87.14(3.11) 84.23(3.42) 5.31(0.68) 8.67(4.52) 5.56(0.68) 4.06(0.60) 5.90(0.75)
FGM 87.07(3.31) 84.16(3.68) 5.43(1.22) 8.86(5.80) 5.74(0.93) 4.39(0.52) 6.10(1.01)
ATC 88.14(1.34) 85.17(1.36) 5.95(0.60) 10.02(2.00) 6.56(1.19) 4.85(0.88) 6.84(0.30)
CL 86.59(2.72) 83.41(3.32) 5.15(0.89) 8.85(1.70) 6.19(0.87) 4.28(0.83) 6.12(0.37)
ER [11] PGD 88.68(0.32) 85.64(0.37) 5.38(0.70) 8.63(0.70) 5.43(0.86) 4.64(0.82) 6.02(0.31)
FGM 88.64(0.55) 85.60(0.68) 5.53(0.23) 9.04(0.74) 5.27(0.42) 4.83(0.87) 6.17(0.21)
ATC 87.17(1.72) 83.83(1.61) 5.10(0.62) 8.58(2.41) 6.32(1.52) 4.79(1.13) 6.20(0.37)
CL 87.19(0.73) 84.05(0.88) 4.96(0.93) 8.99(2.43) 6.25(0.82) 4.83(1.59) 6.26(0.47)
IDBR [25] ATC 88.74(0.87) 85.81(1.40) 5.30(1.34) 9.45(1.99) 6.53(1.30) 4.76(0.48) 6.51(0.51)
CL 87.57(0.80) 84.40(0.90) 4.33(1.18) 8.18(1.43) 5.32(1.24) 4.70(1.21) 5.63(0.64)
REPINA [28] ATC 87.42(0.97) 84.27(1.21) 5.71(0.79) 9.72(3.06) 5.83(0.45) 4.56(1.00) 6.46(0.37)
CL 88.50(0.84) 85.29(1.31) 6.05(1.52) 9.73(4.36) 6.64(1.40) 4.40(1.21) 6.70(1.12)
IRD [29] ATC 88.70(0.77) 85.79(1.00) 5.66(0.70) 9.94(1.37) 6.74(1.10) 5.28(0.62) 6.90(0.45)
CL 87.11(0.97) 83.84(1.19) 4.41(2.18) 8.02(3.92) 5.40(1.37) 4.49(1.65) 5.58(1.14)
TKD [27] ATC 88.86(1.07) 85.96(1.34) 5.58(0.79) 9.66(0.87) 6.37(1.23) 4.81(1.07) 6.60(0.26)
CL 87.02(1.51) 83.70(1.61) 4.58(1.86) 8.42(2.00) 5.33(0.05) 4.51(1.30) 5.71(0.91)
CLF PGD 87.16(2.02) 84.01(2.16) 5.79(0.93) 8.64(1.48) 4.96(0.56) 4.38(0.89) 5.94(0.46)
FGM 86.37(7.63) 83.35(7.12) 5.78(0.93) 8.00(4.16) 4.57(0.52) 4.35(2.65) 5.67(0.70)
ATC 83.85(13.81) 80.51(13.78) 3.90(2.51) 7.33(5.08) 4.06(2.82) 4.22(1.26) 4.88(1.28)
CL 83.94(6.91) 80.31(6.42) 2.79(2.82) 6.21(2.60) 3.61(2.65) 4.84(1.60) 4.36(1.45)
3.2 Debiasing Performance Forgetting of Continuous Debiasing
Initially, we analyze the results of existing methods to determine whether the continuous debiasing of newly emerging attributes will affect the overall fairness of the model, that is, whether the debiasing performance of the model suffers from the catastrophic forgetting problem in the continuous debiasing task. As shown in Table 5, comparing five MTL-based methods (static fairness methods) with five Fine-tune-based methods (continual debiasing methods), it can be found that continual debiasing tasks are prone to be more biased than static debiasing tasks. Fine-tune-based methods can be regarded as static debiasing methods adapted to continuous scenes. These methods perform poorly in continuous scenarios, indicating models struggle with ever-changing biases. The HateDebias dataset closely mirrors real-world bias occurrences, validating that static methods are inadequate under multiple, continual biases. This suggests that in practical applications, more adaptive debiasing methods are needed to improve the overall performance and fairness of models. This underscores the rationale and necessity of the HateDebias dataset proposed in this paper. It closely resembles the occurrence of bias in real-world scenarios, simulating multiple, continuously changing biases, and provides a real environment closer to practical applications.
3.3 Performance of Different Methods on HateDebias
Our experimental findings, detailed in Table 5, indicate that our approach yields substantial improvements in continual debiasing tasks. Static debiasing methods MTL, which can be considered as an upper-bound followed by [25, 29], show significant debiasing effects when dealing with multiple sensitive attributes. In contrast, fine-tuning-based continuous debiasing methods have limited debiasing capabilities, indicating their ineffectiveness in addressing continuous bias. Additionally, the continuous debiasing effect using the ER, REPINA and IRD methods is worse, indicating an inability to effectively reduce bias. These methods are ineffective because the HateDebias dataset closely mirrors real-world bias occurrences, simulating multiple and continuously changing biases. Experimental results show that existing methods perform poorly in these complex, dynamic environments. Finally, the experimental results show that our proposed framework significantly reduces bias in continuous debiasing tasks. Especially with supervised debiasing strategies, the model’s bias is notably reduced at the cost of slight performance trade-offs in downstream tasks. Compared to methods like REPINA, IRD, and TKD, the CLF method demonstrates better debiasing effects and model performance across various debiasing strategies. Experimental results show the proposed method CLF excels in handling multiple and continuously changing biases in complex real-world scenarios, effectively debiasing while maintaining high performance and fairness.
3.4 Performance of Debiasing Preservation Capability
Table 6: The results of debiasing preservation capability.
Method BC
CLF -1.60(2.15)
IDBR -0.61(1.46)
REPINA 0.00(0.01)
TKD -0.82(1.24)
IRD -1.10(1.67)
To verify that our method can avoid catastrophic forgetting, we designed experiments to evaluate the debiasing preservation capability of the model in continuous learning tasks. We propose a Bias Change (BC) metric to assess the model’s debiasing preservation capability:
$BC=\frac{1}{n-1}\sum_{i=1}^{n-1}FPR_{i}^{n}-FPR_{i}^{i},$ (5)
where $FPR_{i}^{i}$ is the false positive rate of the $i$-th bias in the $i$-th stage. Table 6 presents the results of the debiasing preservation capability. These results demonstrate that our proposed CLF method significantly outperforms others in preserving debiasing effects over time. Specifically, the CLF method achieves a BC value of -1.60, indicating substantial bias reduction and showcasing its effectiveness in mitigating bias. In comparison, the REPINA method shows no improvement, while methods like TKD and IDBR exhibit moderate improvements. The poor performance of the comparative methods indirectly reflects the reasonableness of our dataset in terms of capturing the diversity and variability of real-world biases.
Table 7: Experimental results of different sequence lengths.
Sequence Length Method Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ AAB $\downarrow$ (Var.)
2 IDBR 87.51(3.67) 84.52(4.12) 6.37(2.30)
REPINA 87.81(1.91) 84.71(1.68) 6.62(1.99)
IRD 88.10(2.22) 85.10(2.35) 6.89(2.96)
TKD 88.14(1.30) 85.31(1.28) 6.69(1.84)
CLF 85.17(6.79) 81.91(7.13) 5.77(2.23)
3 IDBR 87.96(1.28) 84.90(1.52) 6.09(0.96)
REPINA 88.43(0.54) 85.28(0.84) 6.60(1.12)
IRD 86.75(4.56) 83.75(4.78) 6.00(0.80)
TKD 87.31(1.77) 84.29(2.11) 6.19(0.72)
CLF 85.78(2.35) 81.81(3.23) 4.96(1.55)
4 IDBR 87.57(0.80) 84.40(0.90) 5.63(0.64)
REPINA 88.50(0.84) 85.29(1.31) 6.70(1.12)
IRD 87.11(0.97) 83.84(1.19) 5.58(1.14)
TKD 87.02(1.51) 83.70(1.61) 5.71(0.91)
CLF 83.94(6.91) 80.31(6.42) 4.36(1.45)
3.5 Effects of Different Length Settings on Hatedebias
In practical applications, biases come in various types and forms. To gradually make our dataset approximate real-world scenarios, we experiment with different sequence lengths to evaluate the debiasing effects under different length settings. We compare continuous debiasing tasks with sequence lengths of 2, 3 and 4 to analyze how the diversity and variability of the dataset affect model performance and fairness. As shown in Table  7, we choose CL as the debiasing method and the results reveal several key insights: (1) As the sequence length increases, the performance metrics generally show a decreasing trend. This suggests that longer sequences may introduce more complexity, making it challenging for models to maintain high performance. (2) The AAB values decrease with longer sequence lengths, indicating improved fairness. This trend demonstrates that datasets with longer sequences better capture the diversity and variability of real-world biases, helping models to generalize and debias more effectively. (3) Our proposed method CLF exhibits the best continuous debiasing capability across different sequence lengths, indicating that it can effectively cover datasets in various real-world scenarios.
4 Conclusion
In this paper, we focus on the continuous debiasing task, addressing the problem of debiasing in hate speech scenarios with multi-type and continuously varying bias types. We have introduced a dataset named HateDebiase to benchmark continuous debiasing tasks. The HateDebiase dataset contains 23,276 samples with four bias attributes, corresponding to four sub-datasets within HateDebiase. Based on HateDebiase, we have also proposed a continuous-learning-based framework to eliminate emerging biases while preserving model performance in downstream tasks. We evaluate the debiasing performance of some widely used methods on HateDebiase and demonstrate that the continuous-learning-based method is a promising direction for improving continuous debiasing tasks.
Supplement to
“HateDebias: On the Diversity and Variability of Hate Speech Debiasing”
Appendix organization:
[…]
1
Appendix A1 Related Work
In this section, we briefly review the related literature, including datasets of hate speech detection, static debiasing, continuous debiasing, and continual learning.
Datasets of Hata Speech Detection.
Hate speech detection has been a focal point in natural language processing research, leading to the creation of several datasets aimed at improving the detection capabilities of various models. These datasets vary significantly in their sources, annotations, and the types of biases they address, providing a broad spectrum of resources for developing and evaluating hate speech detection algorithms. One of the earlier datasets is ETHOS [6], which is built from YouTube and Reddit comments validated through a crowdsourcing platform. ETHOS is designed to recognize the target of hate speech, focusing primarily on detecting who the hate speech is directed towards. Another significant dataset is the Measuring Hate Speech corpus [7], which aims to identify the target community and classify whether a sentence contains hate speech. This dataset emphasizes the categorization of the hate speech target, adding a layer of granularity to hate speech detection. Recent advancements have seen the development of datasets addressing specific types of biases. For example, RaceBiasHate [8] uses Twitter data annotated for hate speech and abusive language to tackle racial bias. This dataset is crucial for understanding and mitigating racial biases in hate speech detection systems. Similarly, Geo-CulturalBiasHate [9] evaluates cultural bias in hate speech datasets by leveraging two interrelated cultural proxies: language and geography. This dataset underscores the importance of considering geographic and cultural contexts in hate speech detection. The MultilingualBiasHate corpus [10] expands the scope by including different types of hate speech related to age, country, gender, and race. This multilingual approach provides a comprehensive resource for training and evaluating models on a variety of hate speech scenarios across different languages.
Static Debiasing.
In the field of Natural Language Processing (NLP), bias mitigation strategies [31, 32, 33] fall into two primary categories. The first involves training models to reduce the influence of sensitive attributes during fine-tuning [34, 35], using techniques like projection-based [36], adversarial [17], or contrastive objectives [37]. This approach requires data specifically annotated for the sensitive attribute, with the debiasing effectiveness dependent on the external corpus quality and the potential introduction of new biases. The second category, task-agnostic training, uses general corpora to address biases, such as by identifying and removing gender biases from encoded representations [38] or enhancing model fairness through selective data exclusion [39], increased dropout rates [40], or equalizing objectives [41]. Additionally, methods like the introduction of fairness-triggering perturbations [42] and equi-tuning for group equivariance in immutable models [43] have been explored, focusing on minimizing differences between standard and debiased model features.
Continuous Debiasing.
Current research on debiasing word embeddings mainly addresses static biases and stereotypes, with efforts on pre-trained language models targeting known sensitive attributes [44, 41]. However, unforeseen biases cannot be fully anticipated due to societal changes. Google’s ML-fairness-gym framework aims to simulate continuous debiasing in machine learning, analyzing the impact of algorithms on information systems and human behavior, yet it does not extend to unstructured data, such as text [45]. To date, continuous debiasing in text processing remains unexplored.
Continual Learning.
Continual learning involves updating models with new information while retaining previous knowledge, avoiding catastrophic forgetting [46]. This process entails maintaining generic internal representations for reuse across tasks and dynamically adjusting task-specific parts. Effective strategies include memory-based methods that use prior task exemplars [47, 24] or synthetic data [48], and regularization techniques that minimize changes in learned representations under a fixed architecture [49, 50]. For example, Li and Hoiem [51] proposed "learning without forgetting" (LwF), a technique that introduces regularization to align the model’s output on the current data with that of the previously trained model. Knowledge distillation [52, 53], gradient-based [54] and expansion-based approaches [55] also mitigate forgetting, each targeting the preservation and enhancement of model adaptability over time. Furthermore, some research endeavors have shifted attention toward continual-learning-based methodologies as potential solutions to address model fairness concerns. Churamani et al. [56] introduced a pioneering application of domain-incremental learning as an effective approach for mitigating bias, specifically in the context of enhancing the fairness of facial expression recognition systems. However, there is no work focusing on the continuous debiasing task.
Figure A1: An illustration of stastic debiasing task and continuous debiasing task. In the static debiasing task, the emergence of a new bias requires a complete re-debiasing of all existing biases, resulting in an exponential increase in both training time and costs. Conversely, in the continuous debiasing task, the emergence of a new bias requires only the fine-tuning of this specific bias, eliminating the need for re-training.
Appendix A2 Illustration of Continuous Debiasing and Static Debiasing
The static debiasing primarily focused on adjusting specific biases, as illustrated in Fig. A1. Static debiasing requires re-debiasing all biases whenever new ones appear, consuming considerable training time and resources. Such a cost makes its application impractical in real-world environments where new biases emerge continuously. For instance, if a model is initially trained to mitigate racial bias, and then gender bias becomes apparent, the entire model needs to be retrained to address both biases simultaneously. This retraining process can be highly inefficient and resource-intensive, especially as the number and types of biases continue to grow.
In fact, models in ever-changing environments are expected to adjust targetedly and dynamically to correct these continuously emerging biases, which we define as continuous debiasing, as shown at the bottom of Fig. A1. Continuous debiasing allows for incremental updates to the model to address new biases as they arise without requiring a complete overhaul of the existing bias mitigation strategies. This type of task is significantly more efficient and scalable, making it more suitable for real-world applications where the social and cultural context is constantly evolving.
The importance and rationale of dynamic debiasing lie in its ability to maintain the performance and fairness of models over time. As new biases emerge, they can quickly degrade the effectiveness of a model if not addressed promptly. For example, a social media platform using a hate speech detection model might initially focus on racial slurs. Over time, if there is a rise in hate speech targeting a specific gender or sexual orientation, the model’s effectiveness could diminish unless these new biases are dynamically mitigated. In a continuously changing environment, static debiasing methods fall short as they do not adapt to the new forms of biases that emerge. Continuous debiasing, however, ensures that the models are regularly updated, thus upholding the standards of fairness and reducing the risk of perpetuating new biases. Moreover, dynamic debiasing helps preserve the investment in existing models by extending their useful life and maintaining their relevance without the need for complete retraining.
Appendix A3 Pseudocode of Our Proposed Method
The pseudocode of the implementation process of our proposed continual-learning-based debiasing framework is shown in Algorithm 1.
Algorithm 1 CLF
1:A set of training task $\{{T_{i}}^{n}\}$, the corresponding data set $\{{D_{i}}^{n}\}$, the store ratio $\gamma$. Initial pre-trained language model $LM^{0}$.
2:Trained language model encoder $LM^{n}$
3:Load pre-trained language model $LM^{0}$
4:for $i=1,...,n$ do
5:     Calculate $\mathcal{L}_{cls}$ with $LM^{i}$ (Eq. 9) $\triangleright$ Text Classification
6:     Calculate $\mathcal{L}_{d}$ with $LM^{i}$ (Eq. 10) $\triangleright$ Text Debiasing
7:     if $i\geq 2$ then
8:         Calculate $\mathcal{L}_{g}$ and $\mathcal{L}_{s}$ with $LM^{i}$ (Eq. 7-8) $\triangleright$ Bias Information Regularization
9:         Calculate $\mathcal{L}_{task}$ with $LM^{i}$ (Eq. 6) $\triangleright$ Task Identification
10:         $M_{t}$ = []
11:         for $r=1,...,i$ do $\triangleright$ Memory Replay
12:              $M_{t}\leftarrow\gamma|D_{r}|$ (Eq. 11)
13:              $M_{t}\leftarrow M_{t}\cup M_{t}^{r}$
14:         end for
15:     end if
16:     Calculate overall objective $\mathcal{L}$ with $LM^{i}$ (Eq. 12) $\triangleright$ Overall Objective
17:     Update model parameters with $\mathcal{L}$
18:end for
Appendix A4 Continual-learning-based Framework
A4.1 Overview
We present a continual-learning-based framework for continual debiasing tasks, aiming to enhance the model’s performance and fairness by promoting the learning of attribute-irrelevant features to effectively mitigate bias. For a given task, our proposed method is adopted to the two-stage training backbone, in terms of the initial training stage and the memory replay stage, where the replay-based method can effectively alleviate the classifier bias. During the initial training stage, the model performs training solely on the current task data, allowing the learning of new task-related knowledge. Subsequently, in the memory replay stage, a subset consisting of representative samples from the previous task data is selected to update the memory. The model is further retrained using the balanced memory bank, facilitating the retrieval and consolidation of previously acquired knowledge. Specifically, after training on $(t-1)$-th $(t\geq 2)$ task, $\gamma|D_{t-1}|$ examples are sampled from the training set $D_{t-1}$ into the memory buffer $M$, where $0\leq\gamma\leq 1$ is the store ratio. Data from $M$ is then merged with the next training set $D_{t}$ when conducting the learning of task $t$. What’s more, in the process of continual learning, an important aspect arises from the presence of shared knowledge across different debiasing tasks. Hence, the model must constantly adapt to the specific knowledge related to each debiasing task while avoiding forgetting the old knowledge throughout the learning process. To tackle this challenge, we disentangle sensitive attribute information from raw text representation to obtain a generic space and bias-specific space, introducing an bias information regularization technique to allow interaction between two spaces. The regularization approach aims to preserve shared knowledge while enabling the model to effectively incorporate bias-specific knowledge for different debiasing tasks.
A4.2 Bias Information Regularization
The operation of disentangled enables the separation of generic and bias-specific information in the hidden representations, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. The disentangled concept we define allows the model to learn generic information about downstream tasks and specific information about sensitive attributes, respectively. For a given sentence $x$, a pre-trained language model $B(\cdot)$ is employed to derive the hidden representations $h$, encompassing both generic and bias-specific information. Subsequently, we design two disentangled networks, represented by $G(\cdot)$ and $S(\cdot)$, which play a vital role in extracting the generic representation $g$ and sensitive bias-specific representation $s$ from the previously obtained hidden representations $h$. $G(\cdot)$ and $S(\cdot)$ refer to two distinct feed-forward networks. The dimensions of the representations produced by $G(\cdot)$ and $S(\cdot)$ are identical to the hidden representations $h$. This systematic process enables the generic and bias-specific information to disentangle, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. For the dynamic debiasing task of sensitive attributes, we train a classifier by exploiting information from two representations and allowing different representation spaces to be changed to different extents, maintaining the ability of the models when debiasing the sensitive attributes.
Generic Representation
The generic representation $g$ is derived from the hidden representation feed-forward network $G(\cdot)$ that contains information common to various debiasing tasks within a given task sequence to alleviate forgetting old knowledge. When transitioning from one debiasing task to another, it is expected that the generic information remains relatively consistent across different tasks. A generic representation is used to learn the ability of downstream tasks and the ability of bias mitigation tasks.
Bias-specific Representation
Models require specific attribute information for effective performance in debiasing tasks. It is crucial for models to accurately identify the specific type of bias present within each text to facilitate effective debiasing. Consequently, we introduce a straightforward bias-identifier prediction task targeting the bias-specific representation $s$, which entails the classification of a given example into its corresponding task category. By adopting this auxiliary approach, we promote models’ learning of distinct bias-specific information within $s$. The loss function for the bias-identifier predictor, denoted as $f_{task}$, is formulated as follows:
$\mathcal{L}_{task}=E_{(x,z)\in D_{t}\cup M_{t}}\mathcal{L}_{CE}(f_{task}(S(B(x)),z),$ (6)
where $z$ is the corresponding debiasing task id for $x$, $L_{CE}(\cdot)$ is the cross entropy loss function and $B(\cdot)$ is the pre-trained model for text encoding in our framework. The memory buffer, denoted as $M_{t}$, retains samples from the training sets of prior tasks. This auxiliary setup serves as a valuable mechanism for enhancing task-specific learning and contributes to the robustness and adaptability of the model across different tasks.
Representation Regularization
To cope with catastrophic forget during continually debiasing pre-trained language models employing continual learning, we propose a regularization approach encompassing both generic representations $g$ and bias-specific representations $s$. This regularization is strategically applied to different representations on $g$ and $s$, as preserving knowledge in distinct spaces separately serves to encourage positive transfer and mitigate forgetting. Prior to initiating the training of all modules on debiasing task $t$, we calculate the generic representations and bias-specific representations for all sentences within the training set $D_{t}$ of the current task and the memory buffer $M_{t}$ ($M_{t}\subseteq M$). Specifically, the encoder $B^{t-1}(\cdot)$ and two disentangled networks $G^{t-1}(\cdot)$ and $B^{t-1}(\cdot)$, $D_{t-1}$ trained on the preceding task $t-1$ is employed to compute the generic representation as $G^{t-1}(B^{t-1}(x))$ and the bias-specific representation as $S^{t-1}(B^{t-1}(x))$, thereby preserving generic knowledge from the prior models to alleviate forgetting. These computed generic and bias-specific representations are retained with the representations obtained at task $t$ to perform L2 norm regularization, constraining the updated degree of the generic representation space and the bias-specific representation space. Thus, during the learning for training pairs in debiasing task $t$, we impose two regularization losses respectively:
$\mathcal{L}_{g}=E_{x\in D_{t}\cup M_{t}}||G^{t-1}(B^{t-1}(x))-G^{t}(B^{t}(x))||_{2},$ (7)
$\mathcal{L}_{s}=E_{x\in D_{t}\cup M_{t}}||S^{t-1}(B^{t-1}(x))-S^{t}(B^{t}(x))||_{2},$ (8)
where $||\cdot||_{2}$ is the L2 norm regularization operation.
A4.3 Text Classification
When performing the $t$-th debiasing task, we leverage a fusion of the generic representation $g$ and the bias-specific representation $s$ to facilitate text classification. This combination enables the model to effectively capture both the generic and bias-specific information in the downstream task. To optimize model performance, we minimize the cross-entropy loss:
$\mathcal{L}_{cls}=E_{(x,y)\in D_{t}\cup M_{t}}\mathcal{L}_{CE}(f_{cls}(g\oplus s),y)),$ (9)
where $y$ denotes the corresponding class label assigned to input $x$, and $f_{cls}(\cdot)$ represents the class predictor function. The symbol $\oplus$" signifies the concatenation operation applied to the two representations under consideration.
A4.4 Text Debiasing
In the $t$-th debiasing stage, the model is aimed at debiasing the sensitive attribute $A_{t}$. Specifically, for a given sentence $x$, we adopt the given debiaser $H_{t}(\cdot)$ to mitigate sensitive attribute $A_{t}$ in the text. The debiaser $H_{t}(\cdot)$ is designed to be highly flexible and extensible, allowing for modifications as needed. In our experiments, we utilize four debiasing methods as the backbones of the debiaser $H_{t}(\cdot)$, namely FGM [15], PGD [16], ATC [17, 18, 19] and CL [20, 21, 22]. We minimize the debiaser’s loss:
$\mathcal{L}_{d}=E_{(x,y)\in D_{t}\cup M_{t}}\mathcal{L}_{d}(H_{t}(x)).$ (10)
We denote $\mathcal{L}_{d}(\cdot)$ as the loss function associated with the debiaser, responsible for executing the debiasing operation on the given sample.
A4.5 Memory Replay
Since we only store a small number of samples from the training set of the previous debiasing task to achieve balance replay with the additional memory cost and training time, we need to choose the samples carefully so that the memory buffer $M$ is used efficiently. The previous memory replay methods [24, 11] in continual learning solely focus on recalling representative samples to maintain the performance of fine-tuned downstream tasks. However, these samples may not necessarily be biased and thus may not be suitable for attribute-incremental continual learning methods. To address this limitation, We explore strategies to selectively recall specific samples in order to mitigate model bias.
Based on the task’s definition, it becomes apparent that bias emerges due to variations in the model’s performance among distinct groups of samples characterized by a sensitive attribute. Therefore, recalling samples with a low accuracy of labels under the current model for replaying can further alleviate the bias of the model, because when there are a large number of samples with low accuracy in the subset of the sensitive attributes, the model will have a greater bias for a sensitive attribute.
In the $t$-th debiasing stage, before fine-tuning and debiasing the model, we select the corpus $D_{r}$ for stage $r(r\textless t)$. First, the class predictor $f_{cls}(\cdot)$ is used to predict $D_{r}$, and the probability list $P$ corresponding to the true label of each sample is obtained. We set the number of retained samples as $\gamma|D_{r}|$, and obtain the corresponding probability threshold as $\beta$, then the samples retained in stage $r$ are:
$M_{t}^{r}=\{x\in D_{r}|P(x)\textless\beta\}(M_{t}^{r}\subseteq M_{t}).$ (11)
A4.6 Overall Objective
The final loss of our dynamic debiasing model is:
$\mathcal{L}=\mathcal{L}_{cls}+\alpha\cdot\mathcal{L}_{d}+\sigma\cdot(\mathcal{L}_{g}+\mathcal{L}_{s}+\mathcal{L}_{task}).$ (12)
We set the coefficient of the first loss terms to 1 for simplicity and only introduce two loss weights to tune: $\sigma$ and $\alpha$. The loss weight $\alpha$ represents the debiasing degree of the model, while the loss weight $\sigma$ controls the regularization degree of bias information.
A4.7 Generalizability of Our Framework
Our framework for continuous debiasing is designed to be adaptable and extendable to various continuous learning methods. This includes replay strategies, regularization strategies, and other continual learning approaches. The primary advantage of our framework is its flexibility, allowing for the direct and convenient application of existing methods to the task of continuous debiasing. This adaptability ensures that the framework can be effectively employed for ongoing bias mitigation across diverse and evolving datasets, facilitating improved model fairness and robustness over time. As new techniques and strategies are developed, they can be integrated into our framework, providing a robust and evolving solution for dynamic debiasing challenges.
Appendix A5 Debiasing Methods
We utilize four debiasing methods in the continuous debiasing task, namely the unsupervised debiasing methods Fast Gradient Method (FGM) [15] and Projected Gradient Descent (PGD) [16], as well as the supervised debiasing method, Adversarial Training Classifiers (ATC) [17, 18, 19] and Contrastive Learning (CL) [20, 21, 22].
Fast Gradient Method (FGM): One of the most common methods for debiasing models is unsupervised adversarial learning [57]. We follow the adversarial training method of FGM [15] to add perturbation to BERT’s representation module.
Projected Gradient Descent (PGD): PGD [16] is another unsupervised adversarial learning method, which adds an iterative attack that differs from FGM. The method performs multiple iterations, taking small steps each time, and projecting the perturbation within a specified range.
Adversarial Training Classifiers (ATC): Unlike unsupervised adversarial training, supervised adversarial training [17, 18, 19] tend to train classifiers to predict sensitive attributes for adversarial learning, thereby blurring the model’s identification of sensitive attributes.
Contrastive Learning (CL): Supervised contrastive learning [20] is utilized as an intermediate pre-training approach for mitigating biases in contextual representations. The contrastive learning [21, 22] is aimed at pulling identical entailment pairs along opposite sensitive attribute directions closer.
Appendix A6 Baselines
We use seven baselines Fine-tune [23], Experience Replay (ER) [24, 11], IDBR [25], Multi-task Learning (MTL) [26] and Two-level Knowledge Distillation (TKD) [27], Representation Projection Invariance (REPINA) [28] and (Instance-wise Relation Distillation (IRD) [29] to combine debiasing methods to verify the effectiveness of various strategies and our proposed method. Given that our experimental setting involves attribute-incremental continual learning, we transform the attribute-incremental setting into a debiasing-task-incremental setting to enable the application of the task-incremental continual learning method to our dataset.
Fine-tune [23] modifies the parameters of the pre-trained language model sequentially without any augmented strategies and additional loss.
Experience Replay (ER) [24, 11] is a memory-based technique that involves retaining a limited subset of samples from prior tasks and subsequently reintroducing them during training. The method effectively tackles the issue of catastrophic forgetting, ensuring that the model retains and integrates old knowledge in the learning process.
IDBR [25] is a novel approach that incorporates information disentanglement regularization into the encoding process. This technique disentangles the encoding of task-specific information and general information. By jointly considering these distinct encodings during the classification process, IDBR offers a promising strategy for enhancing the model’s discriminative capabilities and overall performance.
Multi-task Learning (MTL) [26] requires the model trained on all tasks simultaneously, which can effectively avoid the problem of catastrophic forgetting and be considered as a static debiasing method since it has access to data from all tasks at the same time.
Two-level Knowledge Distillation (TKD) [27] is a technique designed to improve the continual text classification process by mitigating catastrophic forgetting. This method involves transferring knowledge at two distinct levels: the feature level and the output level. The TKD strategy operates within a teacher-student framework, where the teacher is the model trained on previous tasks, and the student is the current model being trained. By integrating both feature-level and output-level distillation, the method effectively slows down the forgetting process.
Representation Projection Invariance (REPINA) [28] is a novel regularization method designed to mitigate the issue of representation collapse during the fine-tuning of pre-trained language models. REPINA aims to preserve the information content of representations by discouraging undesirable changes. This approach involves applying an invariance loss on a tunable projection of the representation, allowing it to change mildly without losing its expressivity.
Instance-wise Relation Distillation (IRD) [29] is a regularization technique used in continual learning to mitigate the forgetting of previously learned knowledge when the model learns new tasks. Specifically, it measures and minimizes the discrepancy between the representations of the current model and those of the previous model. This helps maintain the stability of the learned knowledge. By doing so, the model can adapt its representations while learning new tasks without significantly deviating from the previously learned knowledge.
Appendix A7 Supplementary Experiments
A7.1 Ablation Results
Table A1: Ablation experiment results. Upon removal of the BIR module, both the performance and fairness metrics of the models deteriorated.
Debiasing Method Method Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ Gender $\downarrow$ Age $\downarrow$ Country $\downarrow$ Ethnicity $\downarrow$ AAB $\downarrow$
ATC CLF 83.85 80.51 3.90 7.33 4.06 4.22 4.88
w/o BIR 85.73 81.66 4.42 8.67 5.17 4.57 5.71
[…]
w/o BIR 83.66 80.31 4.78 7.44 4.81 4.24 5.32
As shown in Table A1, we conduct ablation experiments to verify the effectiveness of BIR. It can be seen that when ablating BID, the AAB values of CLF with ATC and CLF with CL increased by 0.0083 and 0.0096, reflecting that the model is more biased. The result demonstrates that BIR can effectively alleviate the continual biases in the model. Simultaneously, BIR would affect the performance of the model in downstream tasks to a certain extent. When discarding BID, the accuracy of CLF with ATC and CLF with CL increased by 0.0188 and 0.0028.
We further explore the performance and fairness of the model utilizing our designed replay strategy and the ER strategy, respectively. It can be seen that the performance of BIR solely adopting our replay strategy in downstream tasks is not as good as those ER methods. Whereas from the perspective of the AAB metric for evaluating fairness, the replay strategy makes the model more fair. The experimental results show that the BIR strategy and our proposed memory replay strategy can effectively debias the pre-trained language model in the continuous debiasing task.
A7.2 Performance of Different Task Sequences
Table A2: The experimental results of the CLF with CL. The columns and rows of the table are the evaluation metrics and abbreviations for each task sequence respectively. The column located between the dotted and solid lines represents the average of the six series preceding the dotted line.
Sequence Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ Gender $\downarrow$ Age $\downarrow$ Country $\downarrow$ Ethnicity $\downarrow$ AAB $\downarrow$
aceg 81.15 77.98 1.09 3.94 2.98 4.73 3.18
acge 87.80 84.54 4.13 8.12 6.36 5.40 6.00
aecg 86.26 82.89 3.02 7.12 6.08 5.18 5.35
aegc 85.15 81.58 2.27 7.04 4.12 6.19 4.91
agce 84.35 80.07 0.39 4.74 1.66 4.18 2.74
agec 81.00 77.45 0.14 4.66 3.34 3.42 2.89
[…]
caeg 84.89 81.44 4.93 7.53 4.61 6.92 6.00
cage 84.26 80.39 3.83 4.43 0.41 2.79 2.87
ceag 84.42 80.60 5.83 6.04 4.38 5.10 5.34
cega 84.55 81.18 6.61 8.92 4.56 5.58 6.42
cgae 76.89 73.84 1.38 3.53 3.81 4.05 3.19
cgea 84.65 81.17 2.99 8.42 5.45 8.52 6.35
[…]
eacg 85.77 82.24 3.70 6.53 1.59 3.48 3.82
eagc 86.26 83.08 2.94 8.36 6.10 3.45 5.21
ecag 84.73 81.70 2.32 7.09 5.56 5.47 5.11
ecga 86.22 82.33 2.43 6.29 3.29 4.03 4.01
egac 80.31 76.69 0.35 4.28 1.56 5.41 2.90
egca 80.48 77.64 2.06 5.48 3.20 4.06 3.70
[…]
gace 83.78 79.61 1.64 5.15 2.09 4.49 3.34
gaec 85.08 80.90 2.69 5.45 2.18 4.09 3.60
gcae 85.61 81.72 2.86 6.45 4.21 4.04 4.39
gcea 86.21 81.79 4.51 8.83 3.43 5.49 5.56
geac 79.62 75.80 3.74 4.84 1.80 4.25 3.66
geca 85.08 80.76 1.22 5.83 3.82 5.78 4.16
[…]
Avg 83.94 80.31 2.79 6.21 3.61 4.84 4.36
We further explore the impact of the debiasing task order on the performance and fairness of the model. The experiment results of 24 task sequences for the CLF with CL model are shown in Table A2. When the debiasing order is “agce”, the average bias of the four sensitive attributes in the model is the smallest, with an ABB value of 0.0274, while concerning debiasing order “cgea”, the ABB value reaches 0.0635. We further calculate the average ABB values when the initial debiasing attribute in the debiased sequence is “age”, “country”, “ethnicity”, and “gender”. The results reveal that the selection of the initial debiasing attribute will affect the fairness of the model. When first mitigating the biases related to the sensitive attribute “country”, the model retains more biases. It can also be seen that in this case, the average performance of the model in downstream tasks is the worst. It is possible that the features associated with the country attribute have a great impact on the model’s ability to recognize hate speech. Therefore, when the model initially debiases the country attribute and subsequently forgets its debiasing ability, it amplifies the bias of this attribute.
A7.3 Experiments on Other Pre-trained Models
Table A3: Results of Deberta-v3-base and RoBERTa-base. The columns and rows of the table are the evaluation metrics and the names of each method respectively. Experimental results on different backbones show that our method has obvious continuous debiasing capabilities and is better than existing methods.
Debiasing Method Backbone Method Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ Gender $\downarrow$ Age $\downarrow$ Country $\downarrow$ Ethnicity $\downarrow$ AAB $\downarrow$
CL Deberta -v3-base MTL 87.94 84.90 06.01 10.07 4.91 4.97 6.49
Fine-tune 88.86 85.77 6.23 11.04 5.88 5.16 7.08
IDBR 86.15 82.88 5.63 8.98 4.37 5.46 6.11
CLF 83.95 80.85 5.35 9.24 4.56 4.54 5.92
RoBERTa -base MTL 87.53 84.25 4.98 8.43 4.49 4.45 5.59
Fine-tune 88.16 85.07 5.17 9.07 4.47 4.62 5.83
IDBR 86.22 82.96 4.14 7.32 4.92 3.99 5.09
CLF 85.53 82.13 3.65 7.23 4.42 3.46 4.69
In addition to conducting experiments on the BERT model, we further perform our method with other pre-trained language models Deberta-v3-base [58] and RoBERTa-base [59]. The experimental results are shown in the Table A3. Experimental results verify the efficacy of our method in mitigating biases across various large-scale pre-trained models. This further validates the versatility and applicability of our approach in addressing bias within the realm of natural language processing models, underscoring its potential utility in enhancing the fairness and impartiality of machine learning applications.
A7.4 Parameter Search Experiments
Table A4: The experimental results of parameter exploration. Unlisted values are identical to those of the base model.
Parameters Performance Fairness
$\gamma$ $\alpha$ $\sigma$ Acc $\uparrow$ F1-macro $\uparrow$ Gender $\downarrow$ Age $\downarrow$ Country $\downarrow$ Ethnicity $\downarrow$ AAB $\downarrow$
base 0.1 0.1 0.1 83.94 80.31 2.79 6.21 3.61 4.84 4.36
(a) 0.05 84.20 80.60 2.90 6.15 4.31 5.00 4.59
0.15 85.12 81.74 3.90 7.58 4.43 4.68 5.15
[…]
0.15 85.25 82.00 3.51 7.53 4.66 4.53 5.06
[…]
0.15 83.92 80.32 2.55 6.27 3.98 4.94 4.43
[…]
0.05 0.05 85.09 81.30 3.06 6.83 4.33 4.53 4.69
0.05 0.05 0.15 85.65 82.46 3.77 7.77 5.36 4.70 5.40
[…]
0.05 0.05 85.36 81.81 4.19 7.76 4.53 5.06 5.38
0.15 0.05 0.05 82.85 79.54 3.60 6.99 4.39 3.98 4.74
[…]
0.05 0.05 84.87 81.29 3.54 6.66 4.31 4.83 4.83
0.05 0.15 0.05 84.82 81.50 4.08 7.75 4.92 4.61 5.34
[…]
0.15 0.15 84.53 81.48 4.02 8.04 5.08 4.11 5.31
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
[…]
0.15 0.15 84.94 81.57 3.83 7.24 4.04 4.81 4.98
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
[…]
0.15 0.15 86.15 82.84 3.99 8.33 5.27 4.75 5.58
0.15 0.15 0.15 85.87 82.70 4.60 8.05 5.14 4.15 5.49
Table A5: The detailed experimental results of debiasing preservation capability. The columns and rows of the table are the methods and abbreviations for each task sequence respectively.
Sequence CLF REPINA IDBR DKT RID
aceg -3.22 -0.01 -1.55 -1.27 -1.66
acge -0.64 0.05 0.26 0.05 -2.70
aecg -0.52 0.03 -1.41 -1.45 -2.46
aegc -1.91 0.00 -1.68 -0.81 -0.67
agce -3.03 0.03 -2.55 -0.84 -1.63
agec -4.18 0.00 -0.15 0.23 0.42
caeg -0.23 -0.08 -0.72 -2.94 -1.58
cage -3.07 0.04 -1.57 -1.67 -2.46
ceag -1.85 -0.03 -0.57 -1.33 -2.00
cega -1.45 0.00 0.09 0.76 -1.82
cgae -3.29 0.06 0.64 -1.43 -0.41
cgea -0.43 0.02 0.17 0.31 0.47
eacg -2.25 -0.06 -1.68 -2.81 -3.03
eagc -1.18 -0.03 -3.14 -2.17 -2.35
ecag 0.05 -0.04 -0.72 -1.40 -1.12
ecga -1.28 0.00 -0.67 -1.11 -0.97
egac -4.64 -0.03 -2.35 -1.41 0.76
egca -3.17 -0.05 -0.51 -0.36 -1.01
gace -1.52 0.05 -0.86 -1.70 -3.40
gaec -0.40 0.02 0.24 0.20 0.31
gcae -0.36 0.07 0.79 -0.79 0.06
gcea 0.99 -0.02 0.71 0.27 -0.40
geac -0.35 0.03 1.77 0.60 0.02
geca -0.52 -0.01 0.72 1.32 1.22
Avg -1.60 0.00 -0.61 -0.82 -1.10
We conducted exploratory experiments on different parameters, which are shown in Table A4. Each parameter has a unique influence on the model’s performance and fairness. Isolating the parameters helps to understand their individual effects, but the best setting would likely depend on the specific performance and fairness goals of the study or the application in which the model is to be used:
Impact of $\gamma$ Change: Experiment (a), (e) and (i) show how varying $\gamma$ influences outcomes. With $\alpha$ and $\sigma$ fixed at 0.1, increasing $\gamma$ (from base to experiment a) reduces performance slightly but improves fairness in gender. With $\alpha$ and $\sigma$ at 0.05, increasing $\gamma$ (experiment e) seems to balance both performance and fairness reasonably well. Experiment (i) also increases $\gamma$, but with $\alpha$ and $\sigma$ at 0.15, showing improvements in performance and fairness, except for age and country.
Impact of $\alpha$ Change: Experiment (b), (f) and (h) all tweak $\alpha$. In experiment (b) with $\gamma$ and $\sigma$ at 0.1, decreasing $\alpha$ slightly decreases performance metrics but improves fairness. With $\gamma$ and $\sigma$ at 0.05, the change in $\alpha$ does not significantly impact performance but seems to worsen fairness in age, country, and ethnicity.
Impact of $\sigma$ Change: Experiment (c), (d) and (g) vary $\sigma$ with $\gamma$ and $\alpha$ at 0.05, 0.1 and 1.05 respectively. In these cases, decreasing $\sigma$ tends to improve fairness in gender and ethnicity but at a slight cost to performance. Experiment (g) increases $\sigma$ with $\gamma$ and $\alpha$ at 0.15, showing an improvement in performance but worsening fairness metrics, notably in gender, country, and ethnicity.
Optimal Settings: Experiment (base) appears to find a middle ground, where increasing $\gamma$ to 0.1 with $\alpha$ and $\sigma$ also at 0.1 offers a balance of improved performance and fairness.
A7.5 Detail of Debiasing Preservation Capability Results
The debiasing preservation capability of different methods is critical to ensure that debiasing performance is maintained across successive tasks without significant degradation. We chose CL as the debiasing method and the results of various methods in preserving debiasing effects over training are presented in Table A5. The results of the debiasing preservation capability evaluation highlight the superior performance of the CLF method in maintaining debiasing effects over time. This method’s unique approach of integrating contrastive learning and continual learning frameworks helps in effectively mitigating biases while preserving overall model performance. Other methods like REPINA show no improvement in bias preservation, while methods like IDBR and TKD exhibit moderate improvements. The significant improvements observed with the CLF method suggest that it is a viable solution for continuous debiasing tasks, promoting fairness and reducing biases in pre-trained language models in the long term.
A7.6 Statistical Significance Test
Table A6: Results of Wilcoxon signed-rank test significance analysis. A p value below 0.05 signifies that our method demonstrates statistically significant improvement over the comparison method.
Debiasing Method Method Compare Method p
ATC CLF Fine-tune 3.93e-06
Replay 6.00e-02
IDBR 6.56e-06
CL CLF Fine-tune 6.04e-02
Replay 5.13e-06
IDBR 1.75e-04
We have conducted a statistical significance test to validate the significance of the experiment results. We performed a Wilcoxon signed-rank test significance analysis [60] on the experimental results of our model CLF (on 24 different debiasing sequences) as well as the experimental results of other continual debiasing methods. We chose ATC and CL as the debiasing methods for significance analysis. The experimental results are shown in Table A6. The p values of the significance analysis of our proposed model and the existing model are shown in the table. Experimental results show that our improvement is significant.